Space probes communicate with controllers on Earth using radio waves. Suppose a probe has a radiotransmitter that operates at a frequency of 645.0 MHz when at rest relative to the observer. This probe is now wandering the solar system. The signal you pick up from the probe is shifted 7.831E+4 Hz higher in frequency than the rest frequency given above. How fast is the probe traveling and is it traveling towards or away from Earth? (assume it is either traveling directly towards or directly away from Earth. You must get boths answers correct at the same time.)
Answer #1

Solution:

Using doppler shift ,

As f' < f \Rightarrow probe is moving away from earth at a speed v,

Using the relation,

f' = (v/c) f Here c = velocity of radio waves,

7.831*104 *3*108 = v 645*106

v = 7.831*104 *3*108/645*106

= 3.6*104 m/s Away.

I hope you understood the problem and got your answers, If yes rate me!! or else comment for a better solutions